# Blind and Topological Interference Managements for Bistatic Integrated Sensing and Communication

### v4

@@ V Achievable schemes for the ISAC Model with Heterogeneous Connectivity
This section provides the proposed ISAC schemes based on the TIM strategy for the bistatic ISAC system with heterogeneous connectivity.
@@
V-A Proof of Theorem 4-a: Topology 1 with Non-cooperative Transmitters
Considering the bistatic ISAC systems with heterogeneous connectivity and non-cooperative transmitters, we first describe the main idea of the proposed TIM scheme for the Topology 1 through the following example. In the bistatic ISAC systems with heterogeneous connectivity, all channels do not change during $t_{0}$ time slots; so for the ease of notation, we remove the time indices from channel coefficients.
Example 2 (Topology 1 with non-cooperative transmitters).
Consider the bistatic ISAC system with partially connected interference channel including $6$ single-antenna transmitters, $5$ single-antenna receivers and one sensor, as illustrated in Fig. 6. Each receiver or sensor is disconnected from the previous $U=1$ and next $D=2$ transmitters, and connected to other $3$ transmitters; the topology belongs to Topology 1 from the $(6,1,2)$-neighboring antidotes network described in Definition 2. We have
@@ V-A Proof of Theorem 4-a: Topology 1 with Non-cooperative Transmitters
For this system, the tradeoff points $(\text{sDoF},\text{cDoF})=(1,0)$, $(\text{sDoF},\text{cDoF})=(0,2)$ can be achieved by the sensing-only and communication-only schemes, respectively. We then propose a TIM ISAC scheme, achieving $(\text{sDoF},\text{cDoF})=(1/5,2)$.
Let $t_{0}=15$. During these $15$ time slots, the sensor will obtain $3$ effective observations, through the sensing signals ${\mathbf{x}}_{1}^{[0]},{\mathbf{x}}_{2}^{[0]},{\mathbf{x}}_{3}^{[0]}\in\mathbb{C}^{3\times 1}$, satisfying $[{\mathbf{x}}_{1}^{[0]},{\mathbf{x}}_{2}^{[0]},{\mathbf{x}}_{3}^{[0]}][{\mathbf{x}}_{1}^{[0]},{\mathbf{x}}_{2}^{[0]},{\mathbf{x}}_{3}^{[0]}]^{H}=I$. In each period of $5$ time slots, we let each receiver recover $2$ communication messages, and let the sensor obtain one effective observation. In the following, we illustrate our proposed scheme for the first $5$ time slots, where each receiver $j\in\{1,\ldots,5\}$ should recover $2$ communication messages $W_{1}^{[j]},W_{2}^{[j]}\in\mathbb{C}$.
Denote the transmission signals of transmitter $j$ by ${\mathbf{x}}^{[j]}=[x^{[j]}(1),\ldots,x^{[j]}(5)]^{T}\in\mathbb{C}^{5\times 1}$. Then for the transmitters connected to the sensor, we design
@@
The dimension of the designed signals is $2$, and the dimension of the interference is aligned to $3$. By the linearly independence of ${\mathbf{v}}^{[1]},{\mathbf{v}}^{[2]},{\mathbf{v}}^{[4]},{\mathbf{v}}^{[5]},{\mathbf{v}}^{[6]}$, receiver $1$ can recover $W_{1}^{[1]},W_{2}^{[1]}$ from the first $5$ time slots. Similarly, each receiver can recover $2$ communication messages from the first $5$ time slots; thus the communication degree of freedom is $2$.
@@
Note that $[{\mathbf{v}}^{[2]},{\mathbf{v}}^{[3]},{\mathbf{v}}^{[6]}]$ with dimension $5\times 3$ has $2$ linearly independent left null vectors with high probability; let ${\mathbf{v}}_{0}\in\mathbb{C}^{1\times 5}$ be a non-zero left null vector. Thus we have
${\mathbf{v}}_{0}{\mathbf{y}}^{[6]}={\mathbf{v}}_{0}{\mathbf{v}}^{[3]}H^{[63]}W_{1}^{[3]}+{\mathbf{v}}_{0}{\mathbf{v}}^{[4]}\left(H^{[63]}W_{2}^{[3]}+H^{[64]}W_{1}^{[4]}\right)$
$+{\mathbf{v}}_{0}{\mathbf{v}}^{[5]}H^{[64]}W_{2}^{[4]}+{\mathbf{v}}_{0}H^{[63]}{\mathbf{x}}_{s}^{[3]}+{\mathbf{v}}_{0}H^{[64]}{\mathbf{x}}_{s}^{[4]}$
$+{\mathbf{v}}_{0}H^{[66]}{\mathbf{x}}_{s}^{[6]}+{\mathbf{v}}_{0}{\mathbf{z}}^{[6]}$
$=\begin{bmatrix}H^{[63]}&H^{[64]}&H^{[66]}\end{bmatrix}\begin{bmatrix}{\mathbf{v}}_{0}{\mathbf{x}}_{s}^{[3]}\\ {\mathbf{v}}_{0}{\mathbf{x}}_{s}^{[4]}\\ {\mathbf{v}}_{0}{\mathbf{x}}_{s}^{[6]}\end{bmatrix}+z^{\prime[6]},$
where $z^{\prime[6]}={\mathbf{v}}_{0}{\mathbf{z}}^{[6]}$. Hence, the sensor can eliminate the interference from communication messages, and obtain an effective observation from the first $5$ time slots.
We then fix the values of ${\mathbf{x}}_{s}^{[3]},{\mathbf{x}}_{s}^{[4]},{\mathbf{x}}_{s}^{[6]}$. Recall that $x_{1}^{[0]},x_{2}^{[0]},x_{3}^{[0]}$ has been selected before, satisfying $[{\mathbf{x}}_{1}^{[0]},{\mathbf{x}}_{2}^{[0]},{\mathbf{x}}_{3}^{[0]}][{\mathbf{x}}_{1}^{[0]},{\mathbf{x}}_{2}^{[0]},{\mathbf{x}}_{3}^{[0]}]^{H}=I$. Now we let $x_{1}^{[0]}=[{\mathbf{v}}_{0}{\mathbf{x}}_{s}^{[3]},{\mathbf{v}}_{0}{\mathbf{x}}_{s}^{[4]},{\mathbf{v}}_{0}{\mathbf{x}}_{s}^{[6]}]$; in other words, we solve ${\mathbf{x}}_{s}^{[3]}$ with $5$ parameters satisfying ${\mathbf{v}}_{0}{\mathbf{x}}_{s}^{[3]}=x_{1,1}$, where $x_{1,1}$ is the first element of ${\mathbf{x}}_{1}^{[0]}$, and similarly for ${\mathbf{x}}_{s}^{[4]}$ and ${\mathbf{x}}_{s}^{[6]}$. Therefore, the sensing degree of freedom is $1/5$.
We next generalize Example 2 to achieve the tradeoff point $(\text{sDoF},\text{cDoF})=\left(\frac{1}{{\mathsf{K}}-D+U+1},\frac{{\mathsf{K}}(U+1)]}{{\mathsf{K}}-D+U+1}\right)$ in (4). Each receiver or sensor is disconnected from the previous $U$ and next $D\geq U$ transmitters, and connected to other ${\mathsf{K}}-U-D+1>0$ transmitters; thus the topology could be expressed as
@@
Let $t_{0}=({\mathsf{K}}-D-U+1)({\mathsf{K}}-D+U+1)$, during these $({\mathsf{K}}-D-U+1)({\mathsf{K}}-D+U+1)$ time slots, the sensor will obtain ${\mathsf{K}}-D-U+1$ effective observations, through the sensing signals ${\mathbf{x}}_{1}^{[0]},{\mathbf{x}}_{2}^{[0]},...,{\mathbf{x}}_{{\mathsf{K}}-D-U+1}^{[0]}\in{\mathbb{C}}^{({\mathsf{K}}-D-U+1)\times 1}$, satisfying $[{\mathbf{x}}_{1}^{[0]},{\mathbf{x}}_{2}^{[0]},...,{\mathbf{x}}_{{\mathsf{K}}-D-U+1}^{[0]}][{\mathbf{x}}_{1}^{[0]},{\mathbf{x}}_{2}^{[0]},...,{\mathbf{x}}_{{\mathsf{K}}-D-U+1}^{[0]}]^{H}=I$. In each period of $({\mathsf{K}}-D+U+1)$ time slots, we let each receiver decode $U+1$ communication messages, and let the sensor obtain one effective observation. In the next, we will illustrate our proposed scheme for the first ${\mathsf{K}}-D+U+1$ time slots, where each receiver $k\in\{1,2,...,{\mathsf{K}}\}$ should recover $U+1$ messages $W_{1}^{[k]},\ldots,W_{U+1}^{[k]}\in{\mathbb{C}}$.
Denote the transmitted signals of transmitter $j\in\{1,\ldots,{\mathsf{K}}+1\}$ by ${\mathbf{x}}^{[j]}=[x^{[j]}(1),\ldots,x^{[j]}({\mathsf{K}}-D+U+1)]^{T}\in\mathbb{C}^{({\mathsf{K}}-D+U+1)\times 1}$. For the transmitters connected to the sensor, we design based on the cyclic coding,
${\mathbf{x}}^{[i]}\negmedspace=\negmedspace{\mathbf{v}}^{[i]}W_{1}^{[i]}\negmedspace+\negmedspace{\mathbf{v}}^{[<i+1>_{{\mathsf{K}}+1}]}W_{2}^{[i]}\negmedspace+\negmedspace\ldots\negmedspace+\negmedspace{\mathbf{v}}^{[<i+U>_{{\mathsf{K}}+1}]}W_{U+1}^{[i]}\negmedspace+\negmedspace{\mathbf{x}}_{s}^{[i]},$
@@
for each $i\in{\mathcal{R}}_{s}\setminus\{{\mathsf{K}}+1\}$, where ${\mathbf{x}}_{s}^{[j]},j\in\mathcal{R}_{s}$ are the sensing signals to be determined later. For the transmitters not connected to the sensor, design based on the cyclic coding,
${\mathbf{x}}^{[i]}={\mathbf{v}}^{[i]}W_{1}^{[i]}+{\mathbf{v}}^{[<i+1>_{{\mathsf{K}}+1}]}W_{2}^{[i]}+\ldots+{\mathbf{v}}^{[<i+U>_{{\mathsf{K}}+1}]}W_{U+1}^{[i]},$
for each $i\in\{1,\ldots,{\mathsf{K}}+1\}\setminus{\mathcal{R}}_{s}$. Note that ${\mathbf{v}}^{[1]},...,{\mathbf{v}}^{[U+{\mathsf{K}}]}\in{\mathbb{C}}^{({\mathsf{K}}-D+U+1)\times 1}$ represent ${\mathsf{K}}+1$ random i.i.d vectors, where any ${\mathsf{K}}-D+U+1$ of them are linearly independent. For each $k\in\{1,\ldots,{\mathsf{K}}\}$, by removing the dedicated sensing signals from the received signals, the $k$-th receiver obtains
@@
$=\underbrace{H^{[kk]}({\mathbf{v}}^{[k]}W_{1}^{[k]}\negmedspace+\negmedspace\ldots\negmedspace+\negmedspace{\mathbf{v}}^{[<k+U>_{{\mathsf{K}}+1}]}W_{U+1}^{[k]})}_{\text{desire signals}}$
$+\underbrace{\sum^{k+{\mathsf{K}}}_{i=k+D+1}{\mathbf{v}}^{[<i>_{{\mathsf{K}}+1}]}\sum^{min({\mathsf{K}}-U-D,i)}_{s=max(1,i-U)}H^{[ks]}W_{i-s+1}^{[s]}}_{\text{interference signals}}+{\mathbf{z}}^{[k]},$
where $\max(a,b)$ is the larger value (i.e., $a$ if $a\geq b$, otherwise $b$) and $\min(a,b)$ is the smaller value (i.e., $a$ if $a\leq b$, otherwise $b$). Therefore, the dimension of the interference signals is aligned to ${\mathsf{K}}-D$, by the linearly independece of ${\mathsf{K}}-D$ vectors ${\mathbf{v}}^{[<k+D+1>_{{\mathsf{K}}+1}]}\ {\mathbf{v}}^{[<k+D+2>_{{\mathsf{K}}+1}]}\ \ldots{\mathbf{v}}^{[<k+{\mathsf{K}}>_{{\mathsf{K}}+1}]}$, receiver $k$ can recover $U+1$ messages from the first ${\mathsf{K}}-D+U+1$ time slots. Thus, the communication degree of freedom is $\frac{{\mathsf{K}}(U+1)}{{\mathsf{K}}-D+U+1}$.
@@
$=H^{[{\mathsf{K}}+1,{\mathsf{K}}+1]}{\mathbf{x}}_{s}^{[{\mathsf{K}}+1]}\negmedspace+\negmedspace\sum^{{\mathsf{K}}-D}_{i=1}{\mathbf{v}}^{[i]}\sum^{min({\mathsf{K}}-U-D,i)}_{s=max(1,i-U)}H^{{\mathsf{K}}+1,s}W_{i-s+1}^{[s]}$

### v5

@@ IV-B Proof of Theorem 2: MU-MISO
Remark 2 (Connection of the proposed ISAC schemes based on BIA).
In this section, we present a new joint precoding framework tailored for ISAC systems, which explicitly exploits the heterogeneity in coherence times between receivers and sensor. The key characteristics of the three proposed schemes (for the interference channel, MU-MISO and MU-MIMO channels) can be summarized as follows: The key characteristics of the schemes can be summarized as follows:
  • The common point of the three proposed schemes is to leverage the ZF precoding to eliminate interference among receivers (with the CSI of the communication channel) and the BIA strategy to eliminate the interference from the communication messages to the sensor (without CSI of the sensor channel).
  • The primary distinction among the three schemes arises from the antenna configurations and their implications on interference management. In the interference channel with single-antenna transmitters and receivers, spatial beamforming is infeasible due to the insufficient antenna dimension. Conversely, when the transmitter is equipped with multiple antennas, spatial beamforming gains can be exploited, providing additional degrees of freedom. Both MU-MISO and MU-MIMO scenarios share a common design principle: spatial beamforming combined with repeated symbol patterns across time slots creates structured observations that facilitate the separation of communication and sensing signals.
@@ V Achievable schemes for the ISAC Model with Heterogeneous Connectivity
While previous designs rely on heterogeneous coherence times to separate sensing and communication signals, in this section we consider heterogeneous connectivity in infrastructures like cellular networks. Assume that the network topology (i.e., the user-to-base station connectivity) remains relatively stable over time, but the exact channel state information is unknown. As stated in Section III-B, we will consider two topologies, $({\mathsf{K}},U,D)$ neighboring antidotes network and the $({\mathsf{K}},d)$-regular network.
@@
V-A Proof of Theorem 4: $({\mathsf{K}}+1,U,D)$ neighboring antidotes network
Considering the bistatic ISAC systems with heterogeneous connectivity modeled as a $({\mathsf{K}}+1,U,D)$ neighboring antidotes network, we first describe the main idea of the proposed TIM scheme through the following example. In the bistatic ISAC systems with heterogeneous connectivity, all channels do not change during $t_{0}$ time slots; so for the ease of notation, we remove the time indices from channel coefficients.
Example 2 (ISAC system with $(6,1,2)$ neighboring antidotes network).
Consider the bistatic ISAC system with a partially connected interference channel including $6$ single-antenna transmitters, $5$ single-antenna receivers and one sensor, as illustrated in Fig. 6. Each receiver or sensor is disconnected from the previous $U=1$ and next $D=2$ transmitters, and connected to other $3$ transmitters; the topology belongs to the $(6,1,2)$-neighboring antidotes network described in Definition 1. We have
@@ V-A Proof of Theorem 4: $({\mathsf{K}}+1,U,D)$ neighboring antidotes network
For this system, the tradeoff points $(\text{sDoF},\text{cDoF})=(1,0)$, $(\text{sDoF},\text{cDoF})=(0,2)$ can be achieved by the sensing-only and communication-only schemes, respectively. We then propose a TIM ISAC scheme, achieving $(\text{sDoF},\text{cDoF})=(2/5,2)$.
Let $t_{0}=10$. During these $10$ time slots, the sensor will obtain $4$ effective observations through the sensing signals ${\mathbf{x}}_{1}^{[0]},{\mathbf{x}}_{2}^{[0]},{\mathbf{x}}_{3}^{[0]},{\mathbf{x}}_{4}^{[0]},\in\mathbb{C}^{3\times 1}$, satisfying $[{\mathbf{x}}_{1}^{[0]},{\mathbf{x}}_{2}^{[0]},{\mathbf{x}}_{3}^{[0]},{\mathbf{x}}_{4}^{[0]}][{\mathbf{x}}_{1}^{[0]},{\mathbf{x}}_{2}^{[0]},{\mathbf{x}}_{3}^{[0]},{\mathbf{x}}_{4}^{[0]}]^{H}=\mathbf{I}_{3}$. In each period of $5$ time slots, we let each receiver recover $2$ communication message symbols, and let the sensor obtain $2$ effective observation. In the following, we illustrate our proposed scheme for the first $5$ time slots, where each receiver $j\in\{1,\ldots,5\}$ should recover $2$ communication messages $W_{1}^{[j]},W_{2}^{[j]}\in\mathbb{C}$.
Denote the transmission signals of $j$-th transmitter by ${\mathbf{x}}^{[j]}=[x^{[j]}(1),\ldots,x^{[j]}(5)]^{T}\in\mathbb{C}^{5\times 1}$. Then for the transmitters connected to the sensor, we design
@@
The dimension of the designed signals is $2$, and the dimension of the interference is aligned to $3$. By the linear independence of ${\mathbf{v}}^{[1]},{\mathbf{v}}^{[2]},{\mathbf{v}}^{[4]},{\mathbf{v}}^{[5]},{\mathbf{v}}^{[6]}$, receiver $1$ can recover $W_{1}^{[1]},W_{2}^{[1]}$ from the first $5$ time slots. Similarly, each receiver can recover $2$ communication message symbols from the first $5$ time slots; thus the cDoF is $2$.
@@
Note that $[{\mathbf{v}}^{[2]},{\mathbf{v}}^{[3]},{\mathbf{v}}^{[6]}]$ with dimension $5\times 3$ has $2$ linearly independent left null vectors with high probability; let ${\mathbf{v}}_{0,1},{\mathbf{v}}_{0,2}\in\mathbb{C}^{1\times 5}$ be the linearly independent left null vectors. So
${\mathbf{v}}_{0,i}{\mathbf{y}}^{[6]}\negmedspace=\negmedspace{\mathbf{v}}_{0,i}{\mathbf{v}}^{[3]}H^{[63]}W_{1}^{[3]}\negmedspace+\negmedspace{\mathbf{v}}_{0,i}{\mathbf{v}}^{[4]}\left(H^{[63]}W_{2}^{[3]}\negmedspace+\negmedspace H^{[64]}W_{1}^{[4]}\right)$
$+{\mathbf{v}}_{0,i}{\mathbf{v}}^{[5]}H^{[64]}W_{2}^{[4]}+{\mathbf{v}}_{0,i}H^{[63]}{\mathbf{x}}_{s}^{[3]}+{\mathbf{v}}_{0,i}H^{[64]}{\mathbf{x}}_{s}^{[4]}$
$+{\mathbf{v}}_{0,i}H^{[66]}{\mathbf{x}}_{s}^{[6]}+{\mathbf{v}}_{0,i}{\mathbf{z}}^{[6]}$
$=\begin{bmatrix}H^{[63]}&H^{[64]}&H^{[66]}\end{bmatrix}\begin{bmatrix}{\mathbf{v}}_{0,i}{\mathbf{x}}_{s}^{[3]}\\ {\mathbf{v}}_{0,i}{\mathbf{x}}_{s}^{[4]}\\ {\mathbf{v}}_{0,i}{\mathbf{x}}_{s}^{[6]}\end{bmatrix}+z^{\prime[6]},$
where $z^{\prime[6]}={\mathbf{v}}_{0,i}{\mathbf{z}}^{[6]},i\in\{1,2\}$. Hence, the sensor can eliminate the interference from communication messages, and obtain $2$ effective observations from the first $5$ time slots.
We then fix the values of ${\mathbf{x}}_{s}^{[3]},{\mathbf{x}}_{s}^{[4]},{\mathbf{x}}_{s}^{[6]}$. Recall that $x_{1}^{[0]},\dots,x_{4}^{[0]}$ has been selected before. Now we let $x_{1}^{[0]}=[{\mathbf{v}}_{0,1}{\mathbf{x}}_{s}^{[3]},{\mathbf{v}}_{0,1}{\mathbf{x}}_{s}^{[4]},{\mathbf{v}}_{0,1}{\mathbf{x}}_{s}^{[6]}],x_{2}^{[0]}=[{\mathbf{v}}_{0,2}{\mathbf{x}}_{s}^{[3]},{\mathbf{v}}_{0,2}{\mathbf{x}}_{s}^{[4]},{\mathbf{v}}_{0,2}{\mathbf{x}}_{s}^{[6]}]$; in other words, we solve ${\mathbf{x}}_{s}^{[3]}$ with $5$ parameters satisfying ${\mathbf{v}}_{0,i}{\mathbf{x}}_{s}^{[3]}=x_{i,1}$, where $x_{i,1}$ is the first element of ${\mathbf{x}}_{i}^{[0]}$, and similarly for ${\mathbf{x}}_{s}^{[4]}$ and ${\mathbf{x}}_{s}^{[6]}$. Therefore, the sDoF is $2/5$.
We next generalize Example 2 to achieve the tradeoff point $(\text{sDoF},\text{cDoF})=\left(\frac{U+1}{{\mathsf{K}}-D+U+1},\frac{{\mathsf{K}}(U+1)}{{\mathsf{K}}-D+U+1}\right)$ in (5). Each receiver or sensor is disconnected from the previous $U$ and next $D\geq U$ transmitters, and connected to other ${\mathsf{K}}-U-D+1>0$ transmitters; thus the topology could be expressed as
@@
Let $t_{0}=({\mathsf{K}}-D-U+1)\left\lceil\frac{({\mathsf{K}}-D+U+1)}{U+1}\right\rceil$, during these $({\mathsf{K}}-D-U+1)\left\lceil\frac{({\mathsf{K}}-D+U+1)}{U+1}\right\rceil$ time slots, the sensor will obtain $N=(U+1)\left\lceil\frac{({\mathsf{K}}-D+U+1)}{U+1}\right\rceil$ effective observations, through the sensing signals ${\mathbf{x}}_{1}^{[0]},{\mathbf{x}}_{2}^{[0]},...,{\mathbf{x}}_{N}^{[0]}\in{\mathbb{C}}^{({\mathsf{K}}-D-U+1)\times 1}$, satisfying $[{\mathbf{x}}_{1}^{[0]},{\mathbf{x}}_{2}^{[0]},...,{\mathbf{x}}_{N}^{[0]}][{\mathbf{x}}_{1}^{[0]},{\mathbf{x}}_{2}^{[0]},...,{\mathbf{x}}_{N}]^{H}=\textbf{I}_{({\mathsf{K}}-D+U+1)}$. In each period of $({\mathsf{K}}-D+U+1)$ time slots, we let each receiver decode $U+1$ communication message symbols, and let the sensor obtain $U$ effective observations. Next we will illustrate our proposed scheme for the first ${\mathsf{K}}-D+U+1$ time slots, where each receiver $k\in\{1,2,...,{\mathsf{K}}\}$ should recover $U+1$ communication message symbols $W_{1}^{[k]},\ldots,W_{U+1}^{[k]}\in{\mathbb{C}}$.
Denote the transmitted signals of $j$-th transmitter, $j\in\{1,\ldots,{\mathsf{K}}+1\}$ by ${\mathbf{x}}^{[j]}=[x^{[j]}(1),\ldots,x^{[j]}({\mathsf{K}}-D+U+1)]^{T}\in\mathbb{C}^{({\mathsf{K}}-D+U+1)\times 1}$. For each transmitter connected to the sensor (i.e., $i\in{\mathcal{R}}_{s}\setminus\{{\mathsf{K}}+1\}$), we design based on the cyclic coding,
${\mathbf{x}}^{[i]}={\mathbf{v}}^{[i]}W_{1}^{[i]}+{\mathbf{v}}^{[i+1]}W_{2}^{[i]}+\ldots+{\mathbf{v}}^{[i+U]}W_{U+1}^{[i]}+{\mathbf{x}}_{s}^{[i]},$
@@
where ${\mathbf{x}}_{s}^{[j]},j\in\mathcal{R}_{s}$ are the sensing signals to be determined later. For each transmitter not connected to the sensor (i.e., $i\in\{1,\ldots,{\mathsf{K}}+1\}\setminus{\mathcal{R}}_{s}$), we design based on the cyclic coding,
${\mathbf{x}}^{[i]}={\mathbf{v}}^{[i]}W_{1}^{[i]}+{\mathbf{v}}^{[i+1]}W_{2}^{[i]}+\ldots+{\mathbf{v}}^{[i+U]}W_{U+1}^{[i]}.$
Note that ${\mathbf{v}}^{[1]},...,{\mathbf{v}}^{[U+{\mathsf{K}}]}\in{\mathbb{C}}^{{\mathsf{K}}-D+U+1}$ represent $K+U$ randomly i.i.d vectors, where any $K-D+U+1$ of them are linearly independent.
@@
$=\underbrace{H^{[kk]}({\mathbf{v}}^{[k]}W_{1}^{[k]}\negmedspace+\negmedspace{\mathbf{v}}^{[k+1]}W_{2}^{[k]}\negmedspace+\negmedspace\ldots\negmedspace+\negmedspace{\mathbf{v}}^{[<k+U>_{{\mathsf{K}}+1}]}W_{U+1}^{[k]})}_{\text{desire signals}}$
$+\underbrace{\sum^{k+{\mathsf{K}}}_{i=k+D+1}{\mathbf{v}}^{[i]}\sum^{\min({\mathsf{K}}-U-D,i)}_{s=\max(1,i-U)}H^{[ks]}W_{i-s+1}^{[s]}}_{\text{interference signals}}+{\mathbf{z}}^{[k]}.$
Therefore, the dimension of the interference signals is aligned to $K-D$. By the linear independence of the ${\mathsf{K}}-D$ vectors ${\mathbf{v}}^{[<k+D+1>_{{\mathsf{K}}+1}]},{\mathbf{v}}^{[<k+D+2>_{{\mathsf{K}}+1}]},\ldots,{\mathbf{v}}^{[<k+{\mathsf{K}}>_{{\mathsf{K}}+1}]}$, receiver $k$ can recover $U+1$ messages from the first ${\mathsf{K}}-D+U+1$ time slots. Thus, the cDoF is $\frac{{\mathsf{K}}(U+1)}{{\mathsf{K}}-D+U+1}$.
@@
$=H^{[{\mathsf{K}}+1,{\mathsf{K}}+1]}{\mathbf{x}}_{s}^{[{\mathsf{K}}+1]}\negmedspace+\negmedspace\sum^{{\mathsf{K}}-D}_{i=1}{\mathbf{v}}^{[i]}\sum^{\min({\mathsf{K}}-U-D,i)}_{s=\max(1,i-U)}H^{[{\mathsf{K}}+1,s]}W_{i-s+1}^{[s]}$